OpenStudy (anonymous):

(2√(2)-6√(7))(4√(2)+12√(7))=

OpenStudy (anonymous):

\[2\sqrt{2}(4\sqrt{2} + 12\sqrt{7}) -6\sqrt{7}(4\sqrt{2} + 12\sqrt{7})\]

OpenStudy (anonymous):

\[8(\sqrt{2})^{2} + 24\sqrt{14} - 24\sqrt{14} - 72(\sqrt{7})^{2}\]

OpenStudy (anonymous):

\[8*2 -72*7\] 16 - 504 -488

OpenStudy (anonymous):

if u r satisfied with the answer, pls click GOOD ANSWER button for me :))

OpenStudy (anonymous):

Harkirat you are good

OpenStudy (anonymous):

i try to be....... BUT i am human too and hence make mistakes once in a while { :)) PS : As u must have realised, to solve the question, I have used the DISTRIBUTIVE property (a-b)(c+d) = a(c+d) -b(c+d) to expand and the exponential laws\[\sqrt{m}*\sqrt{n}=\sqrt{m*n}\] and \[(\sqrt{m})^{2} = m\]